### v1

# Cosmological Zoom-In Simulation of Odd Radio Circles as Merger-Driven Shocks in Galaxy Groups

Abstract
Context. A new class of distinct radio objects has been recently discovered, which are now commonly referred to as Odd Radio Circles (ORCs). The origin of these features is yet unclear because their peculiar properties are a challenge for our current understanding of astrophysical sources for diffuse radio emission.
Aims. We test the feasibility and limits of major mergers in galaxy groups as a possible formation channel for ORCs.
Methods. By modeling the assembly of a massive galaxy group with a final virial mass of in a magnetohydrodynamic zoom-in simulation with on-the-fly cosmic ray treatment, we are able to derive the X-ray and radio properties of the system self-consistently and compare them to observations.
Results. We show that the X-ray properties for the simulated system are agreeing with characteristics of observed galaxy groups in the regarded mass range, legitimating the comparison between the radio properties of the simulated halo and those of observed ORCs. A major merger between two galaxies in the simulation is triggering a series of strong shocks in the circumgalactic medium, which in unison are forming a ring if the line of sight is perpendicular to the merger axis. The shock is rapidly expanding in radial direction and quickly reaches the virial radius of the halo. This formation channel can hence readily explain the morphology and large extent of ORCs. However, the inferred radio luminosity of these features is lower than for observed counterparts, while the degree of polarization seems to be systematically overpredicted by the simulation.
Conclusions. Fossil cosmic ray populations from AGN and stellar feedback might be necessary to explain the full extent of the radio properties of ORCs, since diffusive shock acceleration was the only source term for non-thermal electrons considered in this work.
Key Words.:
Galaxies: interactions – Galaxies: evolution – intergalactic medium – Radio continuum: galaxies1 Introduction
The current advent of next-generation radio telescopes is delivering a number of interesting surprises to the astrophysics community. Their unprecedented depth and frequency coverage continues to reveal stunning details in various extraterrestrial settings – from galactic sources such as supernova remnants (e.g. Weiler et al. 1986; Dokara et al. 2021; Ball et al. 2023; Do et al. 2024) over to the plentiful extragalactic zoo of objects like star forming disc galaxies (e.g. Martinsson et al. 2016; Krause et al. 2018; Stein et al. 2023; Heesen et al. 2024), AGN-hosting galaxies (e.g. McKean et al. 2021; Venturi et al. 2022; Koribalski et al. 2024a), as well as halos and radio relics of galaxy clusters (e.g. van Weeren et al. 2010; Rajpurohit et al. 2020; Brüggen et al. 2021; Venturi et al. 2022; Cassano et al. 2023; Koribalski et al. 2024c). However, it seems that yet another kind of distinct source needs to be added to the list of possible radio objects. As their name suggests, Odd Radio Circles (ORCs) are peculiar radio features with edge-brightened ring- or shell-like morphology (Norris et al. 2021; Koribalski et al. 2021, 2024b). The spectral indices measured in these features, that can lie between and (assuming that the emission spectrum follows ), suggest that these regions are filled with shock-accelerated cosmic rays (CRs), emitting via synchrotron in radio. When found at low galactic latitudes (e.g., Koribalski et al. 2024c) it can be difficult to identify a host galaxy, leaving the possibility that these radio rings are supernova remnants in or near our Galaxy (Filipović et al. 2022; Omar 2022a; Sarbadhicary et al. 2023). Bonafide ORCs, however, have a central early-type galaxy, strongly suggesting they are of extragalactic origin (Koribalski et al. 2021; Norris et al. 2022, 2025). Since ORCs lack a counterpart in any other wavelength regime, distance and size measurements of ORCs rely on the redshifts of their putative host galaxies. Judging by the elliptical galaxies visible in optical that are situated near the geometrical centers of ORCs, these objects lie at distances of and often have diameters of several hundred kpc, while the stellar masses of the central galaxies lie around . This means that ORCs have enormous sizes compared to their central host, as their extent can even exceed the virial radius of the host system.
Since the nature of ORCs is still under debate, one needs to be careful in controlling the sample of considered objects, in order to ensure an appropriate comparison. The following set of conditions seems to works best as a common denominator: the ORC has to a) display a ring or shell-like morphology with b) at least one elliptical galaxy in its center, where c) the size of the radio ring is large, i.e. comparable to the virial radius of the central galaxy (group). While the neighbouring ORCs 2 & 3 from the original discovery paper by Norris et al. (2021) – who first reported such radio morphologies and have introduced the term ORC – are likely the lobes of a re-started radio galaxy where one lobe has ring morphology, there is a growing list of different ORCs whose origin is yet unclear. It is important to note, that the observed morphologies can depend on the resolution and frequency of the observation facility. While an object may look diffuse in one image, the characteristic ring emission for ORCs can reveal itself in surveys with higher resolution and sensitivity. The following observed objects are falling under the criteria stated above: ORCs 1 & 4 (Norris et al. 2021), ORC 5 (Koribalski et al. 2021), ORC 6 (Dolag et al. 2023; Koribalski et al. 2025), ORC J2223–4834 (Gupta et al. 2022), SAURON (Lochner et al. 2023), Physalis (Koribalski et al. 2024b), Cloverleaf (Bulbul et al. 2024), ORC J1027–4422 (Koribalski et al. 2024c) and ORC J0219–0505 (Norris et al. 2025). There are several ongoing searches for ORCs and similar systems as for example the study by Gupta et al. (2025).
A variety of extragalactic formation scenarios for ORCs have been proposed. Several works have addressed the scenario of a termination shock due to galactic outflows during starburst events (Norris et al. 2021, 2022; Coil et al. 2024). Then there is the possibility that a black hole is responsible for these features, either via binary mergers of supermassive black holes (Koribalski et al. 2021; Norris et al. 2022), tidal disruption of stars by black holes (Omar 2022b), as well as AGN jets, where the ejected radio lobes can adopt a circular appearance under certain projections and physical conditions (Nolting et al. 2023; Fujita et al. 2024; Lin & Yang 2024). In this regard, Shabala et al. (2024) have pointed out that the radio power of a mere AGN jet would decline too quickly to explain the properties of ORCs, and hence concluded that an additional shock running through preexisting AGN bubbles would be necessary in this scenario. Finally, there is the shock-driven formation channel that was first proposed by Dolag et al. (2023) and Koribalski et al. (2024b). Based on cosmological simulations, the suggested concept was that merger-accelerated shocks – i.e. merger shocks of coalescence events in massive galaxies and groups that run into outer accretion shocks – can produce features that very closely resemble ORCs. Such processes have been studied for a while now in relation to galaxy clusters, where merger shocks are most likely source of for radio relics (e.g. Vazza et al. 2009; Ha et al. 2018; Zhang et al. 2019). But in the case of clusters, these features always lie well within the boundaries of the respective halo. Hence, ORCs could be unprecedented evidence for merger shocks in galaxy and group halos, where their extent is pushing beyond the boundaries of the virial radius. However, the simulations by Dolag et al. (2023) predicted radio luminosities that were too low compared to observations, concluding that the relatively low simulated halo mass of could be a major issue, since the total energy budget is underestimated. This problem was also pointed out by Yamasaki et al. (2024), who explored the feasibility of pure accretion shocks as the source for ORCs analytically, concluding that more complex simulations with a larger halo mass of at least are required, since these are the estimated halo masses of the ORC host galaxies.
In the present work, we are addressing the necessity of such a study. We present a first of its kind zoom-in simulation of a galaxy merger in high resolution, where the final group mass is corresponding to observed halos in ORCs. We performed a non-radiative magneto-hydrodynamic simulation with an on-the-fly treatment of spectral CR electrons. First, we summarize the numerical details of the simulation by stating the important features of the code and initial condition in Sect. 2. In Sect. 3, we give a detailed presentation of the simulation results. We outline the time evolution of the merger event and then describe the halo’s X-ray and radio properties, providing inferred mock observation images. Hereby, a particular focus lies on the polarization of the simulated ORC. Similarities and differences with observations are discussed in Sect. 4, before we summarize our results in Sect. 5.
2 Simulation code and setup
The simulation was carried out with the cosmological code OpenGadget3, which is utilizing a smoothed particle magnetohydrodynamics (SPMHD) scheme to model gas dynamics and a Tree-PM solver for the gravity (release paper in prep.). Its infrastructure is based on its predecessor Gadget2 (Springel 2005) but employs major improvements in the gravity solver (Ragagnin et al. 2016) and SPH implementation (Beck et al. 2016b), where also a revised time step criterion is ensuring the stability of the code around strong shocks in a simulation (Pakmor et al. 2012). We use the Wendland C4 function as the kernel to reconstruct the local hydrodynamic properties of the simulated fluid based on the position and smoothing length of the simulated particles (Dehnen & Aly 2012; Donnert et al. 2013). The code includes both a physically motivated conductivity (Jubelgas et al. 2004; Dolag et al. 2004; Petkova & Springel 2009), as well as an artificial conduction and time-dependent viscosity scheme (Price 2012; Beck et al. 2016b), which provides numerical convergence when compared to other hydrodynamic discretization methods.
The code resolves the magnetohydrodynamic equations in Lagrangian form (Dolag et al. 1999; Dolag & Stasyszyn 2009) in the non-ideal regime, i.e. with non-vanishing resistivity (Bonafede et al. 2011). Non-zero contributions of the magnetic field , that can emerge due to numerical truncation errors, are suppressed by the hyperbolic divergence cleaning method described by Tricco et al. (2016), where the implementation into OpenGadget3 will be described by Steinwandel & Price (in prep.). We run the code with an on-the-fly shock finder (Beck et al. 2016a), that is returning the shock’s hydrodynamic quantities, as well as as the Alfvénic Mach number and shock obliquity (Böss et al. 2023b). Finally, we include cosmic ray electrons with the module CRESCENDO, which is a Fokker-Planck solver for spectral cosmic rays (CRs, Böss et al. 2023b). The subgrid model computes and evolves the distribution function of CRs for each resolution element (i.e. SPH particle) at each time step in the simulation. The source for CRs is diffusive shock acceleration (DSA) – which therefore heavily relies on the output of the built-in shock finder – and applies shock obliquity dependent acceleration efficiencies (Pais et al. 2018; Kang 2024). These injected populations can then further evolve in time via adiabatic losses as well as radiative losses, which are inverse Compton scattering and synchrotron emission for CR electrons (for the computation of the synchrotron emissivity, c.f. Appendix 2 by Böss et al. 2023a). The present simulation was performed in a non-radiative regime, which does not include star formation, but only resolves the dark matter and gas evolution.
The initial condition for the simulation is drawn from the COMPASS111http://www.magneticum.org/complements.html#Compass set. It is a suite of zoom-in setups for a diverse selection of halos, from massive clusters down to Milky Way-like galaxies, which were selected from a dark matter-only cosmological simulation with a boxlength of (e.g. Bonafede et al. 2011). We select the dfrogin region, which produces a galaxy group halo with a final virial mass of (Schlachtberger 2014). The gas mass resolution and smoothing length of the simulation are listed in Table 1. The simulation employs the standard WMAP-7 CDM cosmological parameters (Komatsu et al. 2011).
3 Results
Until , the halo undergoes an assembly process that includes various mergers with different galaxies. Shortly after , the main halo experiences a particularly violent event, with a mass ratio of . It is this major merger that is triggering the shock and CR injection with subsequent radio emission that we identify as an ORC in this paper. In the following sections we will analyze the time evolution and properties of the galaxy group around this merger event.
3.1 Gas and shock morphology
Fig. 1 shows the evolution of the system around the time of the major merger event. The rows represent three different output times which are separated by . The left and right halves of the figure display the X-ray surface brightness in the band and the cosmic ray electron energy density, respectively. The X-ray flux was estimated based on contributions from thermal and line emission (assuming ), using the software Smac (Dolag et al. 2005a). We assume a distance of (corresponding to ), as it is the case for the observed ORC Physalis (Koribalski et al. 2024b). We show two projections in each case, which show the system orthogonal and parallel to the merger axis in the left and right panel, respectively. For the rest of this work, we adopt as the time shortly before the merger, visualized in the first row of Fig. 1. Almost directly after impact, two large shock shells emerge from the center, which together form an almost full circle with a diameter of after (third panel in second row of Fig. 1). While surpassing the diameter of ORC Physalis, these dimensions are generally agreeing well with the sizes of other observed ORCs, since Physalis is actually on the lower size end (c.f. Table 3 by Koribalski et al. 2024b). The clear boundary where the X-ray brightness is dropping off is coinciding with the position of the shock front, traced by the cosmic ray electron population in the simulation. The total magnitude of the X-ray flux is generally in agreement with observations of halos in this mass range (Bulbul et al. 2024). Meanwhile the merger shock is further expanding in radial direction, reaching a diameter of about after the merger event, hence far surpassing the virial radius of the system equal to .
3.2 X-ray properties of the circumgalactic medium
First we will focus on the evolution of the hot gas component and and its inferred X-ray properties. Fig. 2 demonstrates them for two different times – shortly before and after the merger. The first column displays the same X-ray maps as in the first two rows of Fig. 1 in the orthogonal projection, while the second and third column show the radial profiles of the mass-weighted median of the electron density and temperature for the gas of both merging halos, respectively. Note that as a filter for the X-ray emitting gas we applied density and temperature cuts of and , respectively. After the first pericenter passage, the gaseous component of the galaxies is effectively shredded, while the dark matter component still maintains two significant overdensities. Therefore we center the profiles on the two potential minima inferred from the underlying dark matter distribution, which mixes on longer timescales. To minimize contamination by the other halo, the radial profiles are only computed in 3D cones with an opening angle of , pointing away from each other. These cones are indicated by the red and blue shaded regions in the X-ray map on the left in Fig. 2, where the respective color is also adopted for both galaxies in the adjacent radial profiles.
Before the merger, at , the density and temperature profiles of both galaxies suggest two relaxed galaxies, where the gas is in hydrostatic equilibrium. The central density, and hence total mass of one halo (blue line) is approximately two times larger than the other (the exact merger mass ratio is ). At – after the initial pericenter passage – the gas properties exhibit considerable differences both in density and temperature. While the initially less massive galaxy (red line) develops a pronounced core, the other displays a significant drop in density at its inner few kpc, corresponding to a peak in temperature for this galaxy. Both halos exhibit a large temperature jump around a radial distance of , after which the temperature in the CGM of both galaxies drops again. This is coinciding with the position of the shock at this moment, also evident by the drop of the X-ray flux visible in the map on the left of Fig. 2. The pericenter passage led to massive stripping of gas from the central regions of both galaxies, visible by the strong decrease in density compared to the picture before. However, one galaxy (red line) still maintains a much more dense core compared to the other one, whose central density is almost an order of magnitude lower.
This has interesting consequences for the X-ray brightness of both galaxies. The gas luminosity due to thermal X-ray emission is only weakly dependent on temperature, while being quite sensitive to the gas density. Since one galaxy here (red line) has a significantly larger central density, it is consequently outshining the second galaxy. X-ray observations of ORC Physalis revealed that the X-ray flux of the system is centered on one of its two merging galaxies, while the other does not display any significant central emission (c.f. Fig. 2 by Koribalski et al. 2024b). Our simulation manages to reproduce this behavior, considering that the densities of the galaxies are overall very low after the merger, which could be the reason why one halo is still detected, while the other is simply too dilute to be visible in X-ray. Forthcoming deeper X-ray data accumulated for Physalis by (Khabibullin et al., in prep) should allow one to test this prediction.
3.3 Mock radio observations
In Fig. 3 we present mock images of the radio flux density due to synchrotron emission at four different frequencies and the resolved spectral indices of the simulated ORC. The orientation of the sightline is orthogonal to the merger axis and we choose the time of best morphological resemblance to observed counterparts ( after the merger event). We shift the object from the simulation redshift of to , in order to be consistent with the distance to the observed ORC Physalis (Koribalski et al. 2024b)222Strictly speaking, this is slightly affecting the inferred radio properties of the system, since we include redshift dependent energy losses of CRs via inverse Compton scattering with CMB photons. The spectra are therefore biased towards higher spectral indices here, but the changes to the results are insignificant.. The frequencies and gaussian smoothing kernel widths are chosen such that they coincide with those of currently operating radio telescopes: (, LOFAR), (, GMRT), (, ASKAP) and (, ASKAP). The respective telescope and observer frequency is written in each panel, while the scale in the upper left panel is indicating a physical size of . Similarly to other current works, our simulation is underestimating the local magnetic field strength since we lack the ultra-high resolution necessary to resolve the turbulent dynamo and hence the main driver for efficient magnetic field amplification (Steinwandel et al. 2020), as well as astrophysical seeding mechanisms (e.g. Beck et al. 2013; Garaldi et al. 2021). In order to still provide realistic mock images of the synchrotron brightness – which in turn is heavily relying on the magnetic field strength – we calculate the synchrotron emissivity by assuming equipartition between turbulent and magnetic pressure. This assumption implies a magnetic field strength of , where and are the local gas density and its turbulent velocity component, respectively (Zhou et al. 2024; Böss et al. 2024, c.f. Sect. 4 for details). Recent simulations indicate that this saturation of the turbulent dynamo is valid in regions with strong shocks, where the generation and amplification of magnetic fields is particularly efficient (Steinwandel et al. 2024; Zhou et al. 2024).
The diffuse structure visible at low frequencies is successively reduced at increasing frequencies. This is due to the strong energy dependence of the cooling time of cosmic rays. Therefore we see high energy cosmic ray electrons predominantly only at shocks, where they are currently accelerated. Electrons at lower energies, however, will maintain their energy longer after being shock-accelerated and hence will continue to emit in radio downstream of the shock. This is the reason for the difference between the two the spectral index maps and in the right column of Fig. 3, which indicate the spectral slope of the radio flux density at low and high frequencies, respectively (). In these images we applied a radial mask, in order to reduce the noise in the outskirts (indicated by the dashed circle in the upper left panel). The spectral slope at low frequencies is generally flatter (with maximum values of about at the shock front), than at high frequencies (about at shock), where the bulk of cosmic ray electrons has already cooled off to lower energies, causing a steepening of the spectrum. This aspect, as well as a comparison to observed counterparts will be described in Sect. 3.4.
3.4 Emission spectrum and total radio power
According to the formation model we are analyzing in this paper, the radio feature that we identify as an ORC is a merger shock around a massive galaxy or galaxy group, that is triggered along the axis of a major merger. In this scenario, the total radio luminosity would be the highest shortly after the merger event, since that would be the time when the shock can inject many CR electron populations efficiently into the relatively dense medium at small radii. If we assume that CRs get accelerated from the thermal pool according to DSA, the powerlaw index of the injected CR’s energy spectrum will depend only on the Mach number . Rewriting in terms of the spectral index of the emission spectrum yields:
In the limit of a strong shock (), the spectral index will be , while weaker shocks will generate steeper spectra (). After the initial central merger, the shock will travel radially outwards with decreasing effective Mach numbers, producing increasingly steep spectra at the front.
In Fig. 4 we present the time evolution of the ORC’s total luminosity produced by synchrotron emission. For this purpose we extracted the particles around the expanding shock front at each time step and calculated their cumulative radio brightness at 50 frequencies equidistant in space according to Eq. A1 by Böss et al. (2023a). Both panels show the total brightness on the y-axis, while the x-axis represents time and frequency in the left and right panel, respectively. On the left, different colors identify varying frequencies according to the colorbar, while the solid black line is highlighting specifically , namely the frequency in which the ORC Physalis (Koribalski et al. 2024b) was observed. The vertical dashed lines indicate the three times at which the simulation was displayed in the rows of Fig. 1 (). Shortly after , the merger triggers the shock – evident by the sudden jump in luminosity – and reaches its maximum by around . Afterwards the total power is steadily decreasing in all frequencies, with the decline being stronger for higher frequencies. This effect becomes even more apparent on the right side of Fig. 4. This panel shows the cumulative emission spectrum of the simulated ORC at four successive times, which are approximately apart, starting from the time of maximum radio brightness at (yellow line). This early-on, we see a relatively flat spectral index of in the lower frequency range. Moving on to later times, the spectrum is rapidly decreasing in total luminosity, with further spectral steepening at higher frequencies. This is due to two effects: while the shock is traveling outwards its effective Mach number is decreasing, which results in a lower acceleration efficiency in the DSA framework implemented in our code. Newly injected CRs therefore display increasingly steep spectra with time (c.f. Eq. 1), while simultaneously the norm of the spectrum is systematically decreasing. The latter feature is also influenced by the decline in local gas density and magnetic field strength with radius, since these circumstances lead to less gas with new CR populations and lower synchrotron emissivity, respectively. For a fixed time step, the steeper slope at high frequencies is caused by radiation losses, since the energy loss time scale of electrons due to Synchrotron emission and inverse Compton scattering is strongly dependent on their energy. This leads to a more rapid energy loss of CR electrons at higher energy bins, which is reflected in the steeper emission spectrum at higher frequencies, since the contribution to the spectral emission of a charged particle in a magnetic field is closely peaked around an energy dependent maximum (Ginzburg 1979).
Taking the spectral index at the earliest time () as a proxy for the overall maximum over time, we can easily estimate the respective Mach number by rearranging Eq. 1 for . This yields that the maximum Mach number of the shock in our simulation is , which is comparable to similar merger events in galaxy clusters. The maximum power in radio at is , which is about three orders of magnitude lower that what is measured at the ORC Physalis (Koribalski et al. 2024b).
3.5 Polarization
Synchrotron radiation of electrons will be polarized with a polarization angle , i.e. the angle between the electric field vector of the associated radiation, and a chosen unit vector in the sky plane, which is defined to be orthogonal to the line of sight. Since the polarization angle is depending only on local properties, the so called polarization fraction will be relatively high in case of a singular, relatively small emitting cosmic ray cloud where the local magnetic field is approximately uniform. But as the number of contributing electrons along the line of sight is increasing, the fractional polarization of the observed radiation will consequently decrease because the direction of the magnetic field is not homogeneous. Therefore, a polarization analysis of radio observations can be a powerful tool in order to dissect the physical processes relevant in ORCs. Since our simulation is self-consistently evolving cosmic ray electrons, we can provide polarization information of the simulated ORC and compare it with observed counterparts for the first time.
The polarization state of electromagnetic radiation at a given frequency is commonly described by the four Stokes parameters and , because they can be inferred directly from observed quantities. Generally, synchrotron radiation is elliptically polarized. But in the limit of ultra-relativistic electrons, the light cone of radiation – which is defined as the solid angle around the velocity vector in which an observer is receiving a significant signal – becomes sufficiently narrow. In this case, the observer will receive only ?flashes? of synchrotron radiation when the electron is moving almost exactly towards the observer. The associated electric field vector is hence only drawing a line in the sky plane (instead of an ellipse), which is equivalent to linear polarization (Ginzburg 1979; Longair 2011). In this limit, is vanishing and only three Stokes parameters remain. Assuming the line of sight is along the -direction, is defined as the total intensity received by the observer integrated along the line of sight:
where is the synchrotron emissivity at a given frequency and position . Meanwhile and , which are actually carrying the polarization information, are defined by
where is the polarized synchrotron emissivity. Note that the polarization angle will be subject to Faraday rotation along the line of sight. The polarization angle therefore needs to be adjusted at each contributing particle along the line of sight according to the traveled distance by
where is the observed wavelength, RM the rotation measure, the local electron number density and the component of the magnetic field vector along the line of sight. The polarization angle is depending only on the magnetic field components in the sky plane and is given by
Finally, the polarization fraction can be obtained via
Generally, two factors will have an influence on the degree of polarization of synchrotron radiation: i) the energy spectrum of the underlying cosmic ray electron population and ii) the turbulence of the magnetic field and hence of the carrying intergalactic medium. In case of a uniform (i.e. non-turbulent) magnetic field, the energy-integrated polarization fraction only depends on the powerlaw index of the CR electron energy distribution, which can be rewritten in terms of the spectral index of the emission spectrum (c.f. Eq. 5.46 by Ginzburg 1979):
The subscript 0 is indicating here that the Stokes parameters from Eq. 7 were integrated over the whole energy range in this case. Although we are regarding only spectral values in this Section, it is still instructive to keep this limit in mind, since it is providing us with a sense of what to expect for a given energy distribution. If the value for the fractional polarization measured in the simulation is significantly lower than that, it will be likely due to turbulence of the magnetic field along the line of sight, since this will result in an increasingly diverse distribution of polarization angles, which can further ?fan out? due to Faraday rotation.
In Fig. 5 we present the polarization fraction of the simulated ORC at a frequency of given by Eq. 7. Each quantity ( and ) was smoothed according to the resolution of ASKAP at (, c.f. Sect. 3.3), where we assume the same distance as to the ORC Physalis at (Koribalski et al. 2024b). The time, scale, orientation and circular mask of the map is the same as in Fig. 3. For the most part, the simulated ORC has a degree of polarization between . We have analysed the rotation measure for the simulation and found that it is negligible outside the halo center, which means that polarization is dominated by the emission at the shock ring (c.f. Fig. 8). Judging by their emission spectrum, CRs are injected quite efficiently in our simulation and yield a spectral index close to the DSA limit of (c.f. Sect. 3.4). The fact that we arrive at a much lower mean polarization degree than the expectancy value given by Eq. 8, is an indicator for the turbulence that is present inside the shocked plasma along the line of sight, as explained above. The sporadic white spots in Fig. 5 of deviantly high degree of polarization in comparison, are sightlines with singular cosmic ray populations, which are therefore not contaminated with additional emissions along the line of sight. This is mostly a numerical artifact, which is also evident by the overall ?popcorn-like? appearance of the map – although we are already working with very high resolution here, the shock finder (which is highly resolution dependent) is occasionally failing to pinpoint the location of the shock smoothly (c.f. Sect. 3.6). Additionally, the injection of CRs is strongly dependent on the evaluated Mach number, as well as the estimated shock obliquity, that is determining the local acceleration efficiency, which can further add uncertainties to the simulated polarization. We point out, however, that this has no significant impact on the the overall value of the polarization fraction across the shock ring. Fig. 5 therefore qualifies for a comparison with observed ORCs, which will be carried out in the Discussion (Sect. 4).
3.6 Shock indicators
The built-in shock finder in the code OpenGadget3 is based on a geometrical on-the-fly scheme, that utilizes pressure gradients to construct the shock normal, along which then the Rankine-Hugoniot jump conditions are evaluated. After masking cold flows and shear flows based on threshold pressure and velocity jumps, the algorithm is finally inferring shock quantities such as the compression ratio and Mach number from approximated hydrodynamic quantities up- and downstream of the shock (Beck et al. 2016a). As mentioned in Sect. 3.5, the accuracy of the shock finder is strongly dependent on the resolution of the simulation. Specifically, the size of the local smoothing kernel of the respective SPH particle is defining the scale on which shocks are resolved. In our case, the mean smoothing scale of the simulated ORC is on the order of at , i.e. when the size and morphology matches best to observed counterparts. For reference, this is twice as big as the smoothing kernel imitating the resolution limit of ASKAP at , indicated by the small circle in the lower left corner of the polarization map in Fig. 5. This is the reason for the appearance of the spot-like features in the inferred mock radio images – in particular in the polarization features – which are encompassing the extent of the SPH particles in these regions.
In order to illustrate the actual extent and width of the shock region, it is instructive to compare the results with an alternative way to determine the shock position. In particular, gaussian filtering of pressure maps has proven to be quite efficient in revealing pressure jumps that indicate a shock in the medium (see also Dolag et al. 2005b; Ragagnin et al. 2017; Sommer et al. 2024). Naturally, this approach is only feasible during post processing of the simulation, since it requires a 2D projection in the correct plane and visual inspection. Nevertheless it is a convenient tool to double check the conclusion of the shock finder. The process is the following: starting from a pressure map , we smooth the image by applying a 2D gaussian kernel function on each pixel of the original image and obtain a blurred variant . Sharp edges in the original map will still be present, but are naturally attenuated in the new image. If we now subtract the two images from each other, features spanning over a larger scale than the smoothing kernel will vanish and only sharp features remain. Lastly, the contribution of each pixel is normalized by the local value of the blurred image, simply to see each edge feature in the resulting map equally strong, no matter the pressure contrast. The final shock map will be therefore given by
In Fig. 6 we present such a shock map of the system which was generated by the method explained above. The time and scale of the simulated ORC is again the same as in the polarization map in Fig. 5. The upper panel is showing the integrated surface density of the gas, while the bottom panel is displaying a shock map in the same scale. The pressure at each pixel was determined via the Compton- parameter, which can be used as a proxy for the integrated electron pressure along the line of sight. Finally, we used a gaussian smoothing kernel length of pixels, which is corresponding to approximately in this case. The shock map in Fig. 6 is revealing the clear edges in pressure, displaying its strikingly circular geometry which was not as clear in the output of the shock finder.
4 Discussion
We have shown in Sect. 3, that merger-accelerated shocks can produce giant circular radio features around massive galaxies and galaxy groups. Their morphology closely resembles those of a recently discovered class of radio objects, so called Odd Radio Circles, also dubbed ORCs (e.g. Norris et al. 2021, 2022; Koribalski et al. 2021, 2024b). It is yet under debate as to what these peculiar structures are, since they have a number of unusual traits. Particularly, they display i) a remarkably circular, edge-brightened morphology, ii) have large sizes comparable to the virial radius of a central elliptical galaxy group and iii) exhibit exceedingly high radio powers, judging by self-similar scaling of their halo mass (c.f. Figure 6 by Koribalski et al. 2024b).
4.1 Comparison to observations and Dolag et al. (2023)
The effectiveness of galaxy mergers in producing circular radio features matching ORCs was first shown in a pilot study by Dolag et al. (2023), where the authors have conducted cosmological zoom-in simulations of a Milky-Way-like halo with a mass of approximately . Since the simulations were performed in a pure hydrodynamics fashion without magnetic fields, they needed to estimate the radio brightness during postprocessing. That was realized by adding CRs to shock regions, where their amount was chosen such that the integrated CR energy was a fixed fraction of the dissipated shock energy, based on the shock finder output at a single timestep. Finally by assuming a constant and homogeneous magnetic field, they could infer the spectral radio flux based on synchrotron emission. Although the morphology and size of ORCs could be reproduced, these simulations predicted radio powers that were far too low compared to what is measured in ORCs. However it was unclear which assumptions specifically have led to this deficiency. In this work, we are closing two main shortcomings of the study, namely by simulating i) a massive enough galaxy group (), comparable to what is found in observed counterparts and ii) using a magneto-hydrodynamic scheme with self-consistent, on-the-fly treatment of spectral CRs.
As discussed in Sect. 3.3, it is necessary to proceed with caution when evaluating the absolute value of the magnetic field in a cosmological simulation. It tends to be underestimated, because technically one needs extremely high resolution in order to resolve the turbulent dynamo process in the intergalactic medium, which is the main driver in amplifying the magnetic field (e.g. Vazza et al. 2014, 2018; Böss et al. 2024; Steinwandel et al. 2024). Since the radio properties of the system are based on synchrotron emission, this can significantly alter the final result. To obtain more realistic predictions, it is therefore appropriate to compare the simulated magnetic field strength with analytic predictions for the system. In Fig. 7, we show the time-dependent radio luminosity due to synchrotron emission at for the simulated ORC. The three vertical lines represent the three times from the visualization in Fig. 1, which are each about apart, starting from shortly before the merger. The colored regions represent three different models for the magnetic field strength, where the lower end is the prediction by the respective model and the upper end is representing a boost by a factor of 10. The orange area is the luminosity profile determined by the simulated magnetic field strength , while the other two were generated by assuming a balance between i) magnetic field pressure and thermal pressure (, blue) and ii) magnetic field pressure and turbulent pressure (, red). Hence the local magnetic field strength is given by
in the two cases respectively, where is the thermal pressure, the mass density and the turbulent velocity. It is defined as , where is the the local velocity, while the local bulk velocity is computed as the kernel weighted mean velocity of the neighboring particles
In order to keep the models overall realistic, the value of the plasma- and the balance factor are scaled for the whole simulation, such that the field strength in the central regions of the merging halos is in the order of several , which is in accordance to what is measured at galaxies in that mass range (e.g. Chyży & Beck 2004; Opher et al. 2009; Nikiel-Wroczyński et al. 2020). As expected, the two models generally produce larger values for the total magnetic field strength than the simulated one, reflected by the systematically larger inferred radio power. The turbulent model is overall the most powerful one, which is particularly visible at later times.
The results from the pilot study by Dolag et al. (2023) introduced in the beginning of this Section are visualized as the horizontal grey band in Fig. 7, where its width is representing the overall extent of the total radio power inferred in postprocessing of their simulation. Compared to their estimate, our simulation is exceeding it by at least three orders of magnitude at moment of peak brightness – already when using the radio power inferred from the simulated magnetic field strength alone. But even taking the boost of stronger field models into account, we are still not reaching the luminosity of observed counterparts indicated by the various dashed-dotted lines in Fig. 7. Consequently, also the surface brightness is underpredicted by the simulation. The ORC Physalis, for example, has a flux density in the order of in the outer radio rings (c.f. bottom panel of Fig. 7), while the simulated counterpart is about three orders of magnitude lower at (cf. Fig. 3). This means that, although we are now resolving a halo in the right mass regime and reproduce the correct X-ray properties (c.f. Sect. 3.2), there is clearly still an aspect in the picture missing, in order to entirely explain the physical nature of these systems.
4.2 Potential role of AGN and stellar feedback
Assuming that there is no other essential process producing significant radio emission other than synchrotron emission, this lack of total radio luminosity would then mean that either there are not enough relativistic CR electrons present, or/and the local magnetic field strength is too low. Since the maximum Mach number of the shock, when the simulated ORC is also the brightest, is already relatively large (, c.f. Sect. 3.4), a stronger shock would not remedy the CR deficiency, because their acceleration efficiency via DSA would not increase substantially. A stronger magnetic field would in principle also give rise to stronger synchrotron emission. The problem is, that several processes are currently underrepresented by our magnetic field models, which are presumed to contribute to the state of the magnetic field in the medium, e.g. Bell amplification in shocks or magnetization of the halo by highly magnetized or laminar outflows. To test the effect of stronger magnetic fields in our simulation, we would need to increase the field strength across the whole system in order to stay self consistent. But this would push the field strength in the central regions of the galaxies beyond actually observed values, and is therefore not an option. Additionally, we need to consider the relative importance of fossil CR populations in the halo, which would be present in case of prior injection events of CRs into the CGM of the merging galaxies. Typically, enough time would have passed such that these old populations would have cooled down to mildly relativistic energies – meaning that they would not be radio-bright anymore – but still not enough time for those populations to join the thermal pool of electrons. While a strong shock is necessary to accelerate charged particles from the Maxwellian distribution, a relatively weak shock would already suffice to push such fossil CRs back to relativistic energies again, effectively boosting the total radio brightness of the whole system (e.g. Caprioli et al. 2018; Kang 2018).
Apart from merger or accretion shocks, stellar outflows and AGN feedback are the other main sources for these fossil populations and it can be assumed that they should play a role in the two galaxies in our simulation. But since it was performed in a non-radiative fashion, meaning that it did not include star formation, it is not possible to make definite statements about the actual contribution of stars and black holes at this point. A reasonable assumption would be that they would raise the amount of relativistic electrons by a factor of 10, which is reflected by the upper end of the inferred luminosity profiles of the simulation in Fig. 7. But evidently, that would not suffice to reach the observed values. Nevertheless, the fractional polarization is actually also pointing towards the importance of fossil CRs in this system. The simulation produces values mainly between (c.f. Sect. 3.5), while measurements of ORC 1 report lower values mostly between in the interior and reaching 30% at most in the ring (Norris et al. 2022). If there are significant amounts of old CR populations present, one would certainly increase the number of emission centers along the line of sight compared to the case where electrons were accelerated from the thermal pool only. But since also the direction of the magnetic field vector is varying along the line of sight – and the polarization angle is depending on that direction – this would naturally diversify the distribution of polarization angles present in the total emission. Therefore, the total polarization fraction along a given line of sight would decrease in the case of additional fossil CRs populations, which is necessary to get into agreement with observations. Thus, in order to give a concluding answer to the question whether fossil CR populations are actually the missing key to explain all radio features of ORCs, one would need to perform the simulation again with star formation and proper feedback implementation including self-consistent CR treatment, which is a challenge to current state-of-the-art cosmological codes and we will refer to in future work.
5 Summary and conclusion
In this article, we have presented a first of its kind zoom-in simulation of a galaxy group performed in a magnetohydronamic setup with spectral cosmic ray treatment on-the-fly. The scientific aim was to test the limits of a strong contender for a formation channel for Odd Radio Circles (ORCs) – a new class of radio objects whose origin is still a puzzle. The idea is that during a major merger event between two galaxies, the immense energy dissipation can release several strong bow shocks (i.e. merger shocks). If the projection is close to the perpendicular direction to the merger axis, these shocks can form a giant ring, which closely resembles the edge-brightened radio features that are now identified as ORCs in observations.
This simulation incorporates two major improvements compared to the pilot study by Dolag et al. (2023), who introduced this formation path: it is i) resolving a 10 times more massive halo with a final virial mass of about , which is now in agreement with the group mass measured in the center of ORCs and ii) determining CR population and local magnetic field direction self-consistently, rather than adding them in postprocessing. We showed that the simulation reproduces the X-ray surface brightness of a galaxy group in that mass range and demonstrated that the peculiar offset between the X-ray and radio-bright centers observed in ORC Physalis (Koribalski et al. 2024b) can be explained by the difference in gas density at the centers of the merging galaxies (Sect. 3.2). By providing mock observations of the radio surface brightness and spectral index based on working frequencies of LOFAR, GMRT and ASKAP, we demonstrated that the morphology and size agrees well with observations (Sect. 3.3). The emission spectrum of the radio shells is indicating a sonic Mach number of , which is comparable to strong shock events in massive galaxy clusters (Sect. 3.4). The simulation produces a fractional polarization degree of across the radio ring (Sect. 3.5). The pressure excess map provided in Sect. 3.6 reveals the sharp shock boundaries responsible for the radio features.
Compared to observations, the simulation is underpredicting the total radio luminosity and overestimating its polarization degree (Sect. 4). Both issues would be alleviated by factoring in additional sources for relativistic cosmic rays apart from DSA, which is the only acceleration mechanism in the present simulation. Such mechanisms would be the injection of cosmic rays via stellar or AGN feedback, which could significantly contribute to the total energy budget and boost the radio brightness when reaccelerated by the merger shock. We will implement a description of these fossil cosmic rays in our future work and investigate whether they could be the missing ingredient for explaining the observed radio power of ORCs.
Acknowledgements.
AI, KD and IK acknowledge support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. LMB is supported by NASA through grant 80NSSC24K0173. The calculations for the hydrodynamical simulations were carried out at the Leibniz Supercomputer Center (LRZ) under the project pn68va. The following software was used for this work: julia (Bezanson et al. 2017), Matplotlib (Hunter 2007), Smac (Dolag et al. 2005a).References
- Ball et al. (2023) Ball, B. D., Kothes, R., Rosolowsky, E., et al. 2023, MNRAS, 524, 1396
- Beck et al. (2016a) Beck, A. M., Dolag, K., & Donnert, J. M. F. 2016a, MNRAS, 458, 2080
- Beck et al. (2013) Beck, A. M., Dolag, K., Lesch, H., & Kronberg, P. P. 2013, MNRAS, 435, 3575
- Beck et al. (2016b) Beck, A. M., Murante, G., Arth, A., et al. 2016b, MNRAS, 455, 2110
- Bezanson et al. (2017) Bezanson, J., Edelman, A., Karpinski, S., & Shah, V. B. 2017, SIAM Review, 59, 65
- Bonafede et al. (2011) Bonafede, A., Dolag, K., Stasyszyn, F., Murante, G., & Borgani, S. 2011, MNRAS, 418, 2234
- Böss et al. (2024) Böss, L. M., Dolag, K., Steinwandel, U. P., et al. 2024, A&A, 692, A232
- Böss et al. (2023a) Böss, L. M., Steinwandel, U. P., & Dolag, K. 2023a, ApJ, 957, L16
- Böss et al. (2023b) Böss, L. M., Steinwandel, U. P., Dolag, K., & Lesch, H. 2023b, MNRAS, 519, 548
- Brüggen et al. (2021) Brüggen, M., Reiprich, T. H., Bulbul, E., et al. 2021, A&A, 647, A3
- Bulbul et al. (2024) Bulbul, E., Zhang, X., Kluge, M., et al. 2024, A&A, 685, L2
- Caprioli et al. (2018) Caprioli, D., Zhang, H., & Spitkovsky, A. 2018, Journal of Plasma Physics, 84, 715840301
- Cassano et al. (2023) Cassano, R., Cuciti, V., Brunetti, G., et al. 2023, A&A, 672, A43
- Chyży & Beck (2004) Chyży, K. T. & Beck, R. 2004, A&A, 417, 541
- Coil et al. (2024) Coil, A. L., Perrotta, S., Rupke, D. S. N., et al. 2024, Nature, 625, 459
- Dehnen & Aly (2012) Dehnen, W. & Aly, H. 2012, MNRAS, 425, 1068
- Do et al. (2024) Do, T., Kothes, R., Hill, A. S., et al. 2024, ApJ, 977, 271
- Dokara et al. (2021) Dokara, R., Brunthaler, A., Menten, K. M., et al. 2021, A&A, 651, A86
- Dolag et al. (1999) Dolag, K., Bartelmann, M., & Lesch, H. 1999, A&A, 348, 351
- Dolag et al. (2023) Dolag, K., Böss, L. M., Koribalski, B. S., Steinwandel, U. P., & Valentini, M. 2023, ApJ, 945, 74
- Dolag et al. (2005a) Dolag, K., Hansen, F. K., Roncarelli, M., & Moscardini, L. 2005a, MNRAS, 363, 29
- Dolag et al. (2004) Dolag, K., Jubelgas, M., Springel, V., Borgani, S., & Rasia, E. 2004, ApJ, 606, L97
- Dolag & Stasyszyn (2009) Dolag, K. & Stasyszyn, F. 2009, MNRAS, 398, 1678
- Dolag et al. (2005b) Dolag, K., Vazza, F., Brunetti, G., & Tormen, G. 2005b, MNRAS, 364, 753
- Donnert et al. (2013) Donnert, J., Dolag, K., Brunetti, G., & Cassano, R. 2013, MNRAS, 429, 3564
- Filipović et al. (2022) Filipović, M. D., Payne, J. L., Alsaberi, R. Z. E., et al. 2022, MNRAS, 512, 265
- Fujita et al. (2024) Fujita, Y., Kawanaka, N., & Inoue, S. 2024, PASJ, 76, 765
- Garaldi et al. (2021) Garaldi, E., Pakmor, R., & Springel, V. 2021, MNRAS, 502, 5726
- Ginzburg (1979) Ginzburg, V. L. 1979, Theoretical physics and astrophysics
- Gupta et al. (2022) Gupta, N., Huynh, M., Norris, R. P., et al. 2022, PASA, 39, e051
- Gupta et al. (2025) Gupta, N., Norris, R. P., Hayder, Z., et al. 2025, arXiv e-prints, arXiv:2506.08439
- Ha et al. (2018) Ha, J.-H., Ryu, D., & Kang, H. 2018, ApJ, 857, 26
- Heesen et al. (2024) Heesen, V., Wiegert, T., Irwin, J., et al. 2024, A&A, 691, A273
- Hunter (2007) Hunter, J. D. 2007, Computing in Science and Engineering, 9, 90
- Jubelgas et al. (2004) Jubelgas, M., Springel, V., & Dolag, K. 2004, MNRAS, 351, 423
- Kang (2018) Kang, H. 2018, Journal of Korean Astronomical Society, 51, 185
- Kang (2024) Kang, H. 2024, Journal of Korean Astronomical Society, 57, 155
- Komatsu et al. (2011) Komatsu, E., Smith, K. M., Dunkley, J., et al. 2011, ApJS, 192, 18
- Koribalski et al. (2025) Koribalski, B. S., Dolag, K., Khabibullin, I., et al. 2025, arXiv e-prints, arXiv:2507.11781
- Koribalski et al. (2024a) Koribalski, B. S., Duchesne, S. W., Lenc, E., et al. 2024a, MNRAS, 533, 608
- Koribalski et al. (2024b) Koribalski, B. S., Khabibullin, I., Dolag, K., et al. 2024b, MNRAS, 532, 3682
- Koribalski et al. (2021) Koribalski, B. S., Norris, R. P., Andernach, H., et al. 2021, MNRAS, 505, L11
- Koribalski et al. (2024c) Koribalski, B. S., Veronica, A., Dolag, K., et al. 2024c, MNRAS, 531, 3357
- Krause et al. (2018) Krause, M., Irwin, J., Wiegert, T., et al. 2018, A&A, 611, A72
- Lin & Yang (2024) Lin, Y.-H. & Yang, H. Y. K. 2024, ApJ, 974, 269
- Lochner et al. (2023) Lochner, M., Rudnick, L., Heywood, I., Knowles, K., & Shabala, S. S. 2023, MNRAS, 520, 1439
- Longair (2011) Longair, M. S. 2011, High Energy Astrophysics
- Martinsson et al. (2016) Martinsson, T. P. K., Verheijen, M. A. W., Bershady, M. A., et al. 2016, A&A, 585, A99
- McKean et al. (2021) McKean, J. P., Luichies, R., Drabent, A., et al. 2021, MNRAS, 505, L36
- Nikiel-Wroczyński et al. (2020) Nikiel-Wroczyński, B., Soida, M., Heald, G., & Urbanik, M. 2020, ApJ, 898, 110
- Nolting et al. (2023) Nolting, C., Ball, J., & Nguyen, T. M. 2023, ApJ, 948, 25
- Norris et al. (2022) Norris, R. P., Collier, J. D., Crocker, R. M., et al. 2022, MNRAS, 513, 1300
- Norris et al. (2021) Norris, R. P., Intema, H. T., Kapińska, A. D., et al. 2021, PASA, 38, e003
- Norris et al. (2025) Norris, R. P., Koribalski, B. S., Hale, C. L., et al. 2025, MNRAS, 537, L42
- Omar (2022a) Omar, A. 2022a, MNRAS, 513, 101
- Omar (2022b) Omar, A. 2022b, MNRAS, 516, L43
- Opher et al. (2009) Opher, M., Bibi, F. A., Toth, G., et al. 2009, Nature, 462, 1036
- Pais et al. (2018) Pais, M., Pfrommer, C., Ehlert, K., & Pakmor, R. 2018, MNRAS, 478, 5278
- Pakmor et al. (2012) Pakmor, R., Edelmann, P., Röpke, F. K., & Hillebrandt, W. 2012, MNRAS, 424, 2222
- Petkova & Springel (2009) Petkova, M. & Springel, V. 2009, MNRAS, 396, 1383
- Price (2012) Price, D. J. 2012, MNRAS, 420, L33
- Ragagnin et al. (2017) Ragagnin, A., Dolag, K., Biffi, V., et al. 2017, Astronomy and Computing, 20, 52
- Ragagnin et al. (2016) Ragagnin, A., Tchipev, N., Bader, M., Dolag, K., & Hammer, N. J. 2016, in Advances in Parallel Computing, 411–420
- Rajpurohit et al. (2020) Rajpurohit, K., Hoeft, M., Vazza, F., et al. 2020, A&A, 636, A30
- Sarbadhicary et al. (2023) Sarbadhicary, S. K., Thompson, T. A., Lopez, L. A., & Mathur, S. 2023, MNRAS, 526, 6214
- Schlachtberger (2014) Schlachtberger. 2014, A Disk-Disk Major Merger Event in a Zoom-Simulation, Master’s thesis, University of Munich
- Shabala et al. (2024) Shabala, S. S., Yates-Jones, P. M., Jerrim, L. A., et al. 2024, PASA, 41, e024
- Sommer et al. (2024) Sommer, J. S., Dolag, K., Böss, L. M., et al. 2024, A&A, 691, A38
- Springel (2005) Springel, V. 2005, MNRAS, 364, 1105
- Stein et al. (2023) Stein, M., Heesen, V., Dettmar, R. J., et al. 2023, A&A, 670, A158
- Steinwandel et al. (2024) Steinwandel, U. P., Dolag, K., Böss, L. M., & Marin-Gilabert, T. 2024, ApJ, 967, 125
- Steinwandel et al. (2020) Steinwandel, U. P., Dolag, K., Lesch, H., et al. 2020, MNRAS, 494, 4393
- Tricco et al. (2016) Tricco, T. S., Price, D. J., & Bate, M. R. 2016, Journal of Computational Physics, 322, 326
- van Weeren et al. (2010) van Weeren, R. J., Röttgering, H. J. A., Brüggen, M., & Hoeft, M. 2010, Science, 330, 347
- Vazza et al. (2014) Vazza, F., Brüggen, M., Gheller, C., & Wang, P. 2014, MNRAS, 445, 3706
- Vazza et al. (2018) Vazza, F., Brunetti, G., Brüggen, M., & Bonafede, A. 2018, MNRAS, 474, 1672
- Vazza et al. (2009) Vazza, F., Brunetti, G., & Gheller, C. 2009, MNRAS, 395, 1333
- Venturi et al. (2022) Venturi, T., Giacintucci, S., Merluzzi, P., et al. 2022, A&A, 660, A81
- Weiler et al. (1986) Weiler, K. W., Sramek, R. A., Panagia, N., van der Hulst, J. M., & Salvati, M. 1986, ApJ, 301, 790
- Yamasaki et al. (2024) Yamasaki, S., Sarkar, K. C., & Li, Z. 2024, MNRAS, 528, 3854
- Zhang et al. (2019) Zhang, C., Churazov, E., Forman, W. R., & Lyskova, N. 2019, MNRAS, 488, 5259
- Zhou et al. (2024) Zhou, M., Zhdankin, V., Kunz, M. W., Loureiro, N. F., & Uzdensky, D. A. 2024, ApJ, 960, 12
Appendix A Rotation measure
Fig. 8 shows the correction to the polarization angle due to Faraday rotation along the line of sight (c.f. Eq. 5) of the simulated halo. We assume a frequency of which corresponds to a wavelength of . The contribution in the central region is clearly dominating, while the outer region at the ORC position is adding insignificant rotation. This is mostly due to the much smaller gas density at these radii. Therefore Faraday rotation of synchrotron emission plays only a minor role in determining the polarization fraction at the shock front (c.f. Sect. 3.5).
Appendix B Entropy features
Applying the same technique as described in Sect. 3.6, we show a projected entropy map of the simulated system in Fig. 9, where the projection and time is again the same as of the pressure map in Fig. 6. As a proxy for the entropy we use
where and are the local temperature and density of the gas respectively. No coherent global structures are visible. Instead, the system displays a random entropy excess pattern, which is indicative of the turbulent motion in action inside the circumgalactic medium. It is mostly driven by the small satellite structures orbiting within the group halo.